OpenStudy (anonymous):

check my work please will give medals if you tell if this is right or not!!! Let f(x)=x+5 and g(x)=x^2. find (g o f) (4) A. 9 <--- my answer B. 81 C. 20 D. 1 2. police can estimate the speed of a vehicle before the brakes are applied using the formula 0.75d=s^2/30.25, where s is the speed in miles per hour and d is the length of the vehicle's skid marks. what was the approximate speed of a vehicle that left skid marks measuring 140 feet? A. about 10 miles per hour <-- my answer B. about 75 miles per hour C. about 15 miles per hour D. about 56 miles per hour

OpenStudy (anonymous):

3. you can model the population of a certain city between 1955-2000 by the radial funciton P(x)=55000square root (x-1945). using this model, in which year was the population of that city 220,000? A. 1949 <--- my answer B. 1961 C. 1953 D. 1957 4.what is the inverse of f if f(x)=3^square root (x-4) A. f^-1(x)=x^3+64 B. f^-1(x)=x^3+4 <--- my answer C. f^-1(x)=4-^3 square root (x) D. f^-1(x)=1/3^ square root (x-4)

OpenStudy (anonymous):

Sorry, question A will be 81.

OpenStudy (isaiah.feynman):

Second answer will be D.

OpenStudy (isaiah.feynman):

Hey @Saeeddiscover do you like Paul Dirac?

OpenStudy (anonymous):

Yes, why?

OpenStudy (anonymous):

so does that mean i did 3 and 4 right ?

OpenStudy (isaiah.feynman):

@rockstargirl17 is the last question \[f(x)=3^{\sqrt{x-4}}\]

OpenStudy (anonymous):

3. 220000 = 55000square root (x-1945) divide both sides by 55000 4 = sqrt(x-1945) square 16 = x - 1945 x= 1961

OpenStudy (anonymous):

Question D doesn't seem to be solved right. I am sorry ....

OpenStudy (anonymous):

Oops! I meant question 4. could you tell me what is the option 3 in the question 4? it doesn't seem to be correct ....

OpenStudy (anonymous):

Hey @rockstargirl17????

OpenStudy (anonymous):

4.what is the inverse of f if f(x)=3^square root (x-4) A. f^-1(x)=x^3+64 B. f^-1(x)=x^3+4 <--- my answer C. f^-1(x)=4-^3 square root (x) D. f^-1(x)=1/3^ square root (x-4)

OpenStudy (anonymous):

@Saeeddiscover

OpenStudy (anonymous):

Option C is somehow given incorrectly. In any case, other options would never be the answer. I think option C can be true here.

OpenStudy (isaiah.feynman):

Look for question 4, if the function is an exponential then its inverse MUST be a logarithmic function! I don't see any logarithms in the options which suggests the answer is not among the options! @Saeeddiscover @rockstargirl17

OpenStudy (anonymous):

You don't need to mix himself. If (a, b) is a point belonging to the function f, so (b, a) must be a point belonging to its inverse. Suppose (a, b) = (3, 5), Is there any function in the list to satisfy (5, 3)?

OpenStudy (anonymous):

Do you think it is an exponential function? @Isaiah.Feynman

OpenStudy (anonymous):

I don't think so. sorry ...

OpenStudy (isaiah.feynman):

Then what is f(x)=3^square root (x-4)?? It can only be

OpenStudy (isaiah.feynman):

\[f(x)=3^{\sqrt{x-4}}\]

OpenStudy (anonymous):

It is a simple function.

OpenStudy (isaiah.feynman):

A simple exponential function you mean? :D

OpenStudy (anonymous):

No. exponential functions are not given in forms like this. an exponential function is defined as the inverse of a logarithmic function f and is written as y = exp(x)

OpenStudy (isaiah.feynman):

Okay. That explains why I get a strange "inverse function" because I have being treating it as an exponential function.